please write clearly on paper and show all steps and circle the correct answer. Thank you!
Use implicit differentiation to the curve ex+2y = x + xy +5. Then find the slope of the curve at the point (-2,1).
Answer #1

Grwen put24 = x they t5. en e 24 = x tay +5. differentiation with respect Implicit to x (x 424) - ( + xy +5) ext2 f (x+2y) =luv) = u nas + vai a ext24 (1729) = 1+ ( x dit ty (29] 57 x24 (172g) = 1 + xy + you)e) (ut24 & 2 entrs (y) =ltryty. -) 228429 y - xy Hy-ex324. 29 -X)Hyrekty Yta entry 2extry axit tla ex+2y de ayl- 2eX utzy62,1) ... 2+ 201) Ita e 26 2+2(1) - (-2) Hl- e Itin 2 eta 2 +2 : (slope = 4